a. Why would afirmthat incurs losses choose to produce rather than shut down?

b. The supply curve for afirmin the short run is the short-run marginal cost curve (above the point of minimum average variable cost). Why is the supply curve in the long runnotthe long-run marginal cost curve (above the pointof minimum average total cost)?

c. In long-run equilibrium, all firms in the industry earn zero economic profit. Why is this true?

4.Given the following demand equation P = 60- 2Q and the cost equation TC(q) = 100+3Q2

For the market of shoes

a.If the enterprises in this market take advantages of perfectly competitive market, find the equilibrium outputs and price.

b.If the seller in this market takes the advantages of pure monopoly, find the equilibrium outputs and price.

c.If the seller in this market takes the advantages ofoligopolistic market, find the equilibrium outputs and price.


1
Expert's answer
2022-03-20T19:08:07-0400

a) Why a firm that incurs losses choose to produce rather than shut down.

If a firms revenue covers all the variable costs incurred but not the total cost then the firm is better off producing in the short run even though it is incurring losses.


b) Long-run supply is not the sum of the existing firms' long-run marginal cost curves. The long-run supply curve is dependent on the number of firms in the market and on how their costs change due to any changes in input costs.


c) In long-run equilibrium, all firms in the industry earn zero economic profit. Why is this true?

This is true because;

  • a positive profit would induce firms to enter, decreasing price and profit, and a negative profit would induce firms to produce less, decreasing price and profit.
  • firms are price takers, maximizing profit by producing where price equals marginal cost
  • barriers to entry and exit prevent firms from earning positive or negative economic profit.

c) For perfectly competitive markets,

the enterprises are price takers

"P= 60-2Q"

"TC= 100+3Q^2"


For price takers

"P = MC =MRC"

"MC= \frac{\Delta TC}{\Delta Q}"

"MC = 6Q"


In equilibrium,

MC= P

"6Q= 60-2Q"

"Q=7.5"


While P is,

"P= 60- 2(7.5)"

"P= 60-15"

"P= 45"


d) For a monopoly equilibrium is the same as in perfectly competitive markets;

"MC = MR"

Total revenue

"P\u00d7Q = 60Q-2Q^2"

"MR = 60- 4Q"


In equilibrium,

"60-4Q = 6Q"

"Q = 6"


Price,

"P= 60 - 2(6)"

"P = 48"